\begin{thebibliography}{99}

\bibitem{BiLe} P.-Y. Bienvenu, T. H. Lê,
Linear and quadratic uniformity of the Möbius function over $\Fq[t]$.
Mathematika 65 (2019), no. 3, 505--529.

\bibitem{Bou13} J. Bourgain, Prescribing the binary digits of primes, Israel J. Math. 194 (2013), no. 2, 935–955.

\bibitem{Bou15} J., Bourgain, Prescribing the binary digits of primes, II, Israel J. Math. 206 (2013), no. 1, 165--182.


\bibitem{DMR19} M. Drmota, C. Mauduit, J. Rivat, Normality along squares, J. Eur. Math. Soc. (JEMS) 21 (2019), no. 2, 507--548.

\bibitem{DMS15} C. Dartyge, C. Mauduit, A. Sárközy, Polynomial values and generators with missing digits in finite fields, Funct. Approx. Comment. Math. 52 (2015), no. 1, 65--74.

\bibitem{DMW21} C. Dartyge, L. Mérai, A. Winterhof, On the distribution of the Rudin-Shapiro function for finite fields, Proc. Amer. Math. Soc. 149 (2021), no. 12, 5013–5023.


\bibitem{DS13} C. Dartyge, A. Sárközy, The sum of digits function in finite fields, Proc. Amer. Math. Soc. 141 (2013), no. 12, 4119--4124.

\bibitem{DES16} R. Dietmann, C. Elsholtz,  I. E. Shparlinski, Prescribing the binary digits of squarefree numbers and quadratic residues, Trans. Amer. Math. Soc. 369 (2017), no. 12, 8369--8388.


\bibitem{Ha16} J. Ha, Irreducible polynomials with several prescribed coefficients. Finite Fields Appl. 40 (2016), 10--25.

\bibitem{LiNi97} R. Lidl, H. Niederreiter, Finite fields, 2nd ed., Encyclopedia of Mathematics and its Applications, vol. 20, Cambridge University Press, Cambridge, 1997.

\bibitem{MW22} M. Makhul, A. Winterhof, Normality of the Thue-Morse function for finite fields along polynomial values. Res. Number Theory 8 (2022), no.3, Paper No. 38, 17pp.


\bibitem{MR09}
C. Mauduit and J. Rivat, La somme des chiffres des carrés, Acta Math. 203 (2009), no. 1, 107--148.


\bibitem{MR10} C. Mauduit,  J. Rivat, Sur un problème de Gelfond: la somme des chiffres des nombres premiers, Ann. of Math. (2) 171 (2010), no. 3, 1591--1646.

\bibitem{M19} S. Mattheus, Trace of products in finite fields from a combinatorial point of view, SIAM J. Discrete Math. 33 (2019), no. 4, 2126--2139.


\bibitem{M16} J. Maynard, Primes with restricted digits, Invent. Math. 217 (2019), no. 1, 127--218.

\bibitem{Mer24+} L. Mérai, Smooth polynomials with several prescribed coefficients, preprint, arXiv: 2307.10280. 

\bibitem{MeWi22} L. Mérai, A. Winterhof, Pseudorandom sequences derived from automatic sequences,  Cryptogr. Commun. 14, 783--815 (2022). 


\bibitem{M18} C. Müllner, The Rudin-Shapiro sequence and similar sequences are normal along squares, Canad. J. Math. 70 (2018), no. 5, 1096--1129.

\bibitem{Po13}  P. Pollack, Irreducible polynomials with several prescribed coefficients, Finite Fields Appl. 22 (2013)
70–78.

\bibitem{P19} S. Porritt, Irreducible polynomials over a finite field with restricted coefficients, Canad. Math. Bull. 62 (2019), no. 2, 429--439.

\bibitem{PoPhD}  S. Porritt, Some topics in the analytic number
theory of polynomials over a finite field, University College London, PhD thesis, 2020.

\bibitem{SW19} Z. Sun,  A. Winterhof, On the maximum order complexity of subsequences of the Thue-Morse and Rudin-Shapiro sequence along squares, Int. J. Comput. Math. Comput. Syst. Theory 4 (2019), no. 1, 30--36.

\bibitem{Sw18} C. Swaenepoel, On the sum of digits of special sequences in finite fields, Monatsh. Math. 187 (2018), no. 4, 705--728.

\bibitem{Sw18a} C. Swaenepoel, Prescribing digits in finite fields, J. Number Theory 189 (2018), 97--114.

\bibitem{Sw18b} C. Swaenepoel, Trace of products in finite fields, Finite Fields Appl. 51 (2018), 93--129.



\bibitem{Sw20} Cathy Swaenepoel, Prime numbers with a positive proportion of preassigned digits, Proc. Lond. Math. Soc. (3) 121 (2020), no. 1, 83--151.
\end{thebibliography}

\documentclass[12pt]{amsart}
\usepackage[utf8]{inputenc}

\usepackage{amsmath,amssymb,amsbsy,amsfonts,amsthm,latexsym,mathabx,amsopn,amstext,amsxtra,euscript,amscd,stmaryrd,mathrsfs,cite,array,mathtools,enumerate,enumitem}

\usepackage{hyperref}
\hypersetup{
    colorlinks=true,
    linkcolor=blue,
    filecolor=magenta,      
    urlcolor=cyan,
    citecolor=magenta,
    pdftitle={Smooth polynomials},
    pdfpagemode=FullScreen,
    }

\newtheorem{theorem}{Theorem}
\newtheorem{lemma}[theorem]{Lemma}

\newcommand{\F}{\mathbb{F}}
\newcommand{\K}{\mathbb{K}}
\newcommand{\ZZ}{\mathbb{Z}}

\newcommand{\CC}{\mathbb{C}}

\newcommand{\Fq}{\mathbb{F}_q}
\newcommand{\e}{\mathbf{e}}
\newcommand{\mand}{ \quad \text{and} \quad}

 \usepackage{todonotes}

\def\cB{{\mathcal B}}

\def\cG{{\mathcal G}}

\def\cM{{\mathcal M}}

\def\cP{{\mathcal P}}

\def\cR{{\mathcal R}}
\def\cS{{\mathcal S}}

\def\cU{{\mathcal U}}
\def\cV{{\mathcal V}}

\def\ba{{\mathbf a}}
\def\bx{{\mathbf x}}
\def\x{{\mathbf x}}
\def\y{{\mathbf y}}
\def\h{{\mathbf h}}

\def\fM{{\mathfrak M}}
\def\fm{{\mathfrak m}}

\newcommand{\rk}{\mathop{\mathrm{rank}}}

\title[Rudin-Shapiro function along irreducible polynomials]{Rudin-Shapiro function along irreducible polynomials over finite fields}
 \author[L.~M{\'e}rai]{L{\'a}szl{\'o} M{\'e}rai}
 \address{Department of Computer Algebra, Eötvös Loránd University,  Pázmány P. sétány 1/c, H-1117, Budapest, Hungary} 
 \email{merai@inf.elte.hu}

\allowdisplaybreaks

\date{\today}

\keywords{irreducible polynomials, finite fields, exponential sums, Rudin-Shapiro function}
\subjclass[2020]{11A63,11T23, 11T30}

\begin{document}

\maketitle

\begin{abstract}
    Let $q$ be an odd prime power and  $\Fq$ be the finite field of $q$ elements. We define the Rudin-Shapiro function $R$ on monic polynomials $f=t^n+f_{n-1}t^{n-1}+\dots + f_0\in\Fq[t]$ over $\Fq$ by 
    $$
R(f)=\sum_{i=1}^{n-1}f_if_{i-1}.
    $$
    We investigate the distribution of the Rudin-Shapiro function along irreducible polynomials. We show that the number of irreducible polynomials $f$ with $R(f)=\gamma$ for any $\gamma\in\Fq$ is asymptotically $q^{n-1}/n$ as $n\rightarrow\infty$.  
\end{abstract}

\section{Introduction}

In recent years, many spectacular results have been obtained on important problems combining some arithmetic properties of integers and some conditions on their digits in a given basis, see for example \cite{Bou13,Bou15,DMR19,MR09,MR10,M16,M18,SW19,Sw20}.
In particular, Mauduit and Rivat \cite{MR10} showed that the \emph{Rudin-Shapiro sequence} along prime numbers is well-distributed.

A natural question is to study analog problems in finite fields, see for example \cite{DMS15,DS13,DES16,DMW21,Ha16,Mer24+,MW22,M19,Po13,P19,Sw18,Sw18b,Sw18a}.
Many of these problems can be solved for finite fields although their analogs for integers are actually out of reach.

In particular, Dartyge and Sárközy \cite{DS13} investigated the distribution of the \emph{Thue-Morse function} in finite fields over polynomial values (see also \cite{DMW21}) and Dartyge, Mérai and Winterhof \cite{DMW21} studied the frequency problem of the \emph{Rudin-Shapiro function} in finite fields over polynomial values. 

Moreover, Porritt \cite{PoPhD} studied the distribution of linear forms of irreducible polynomials as a generalization of the Thue-Morse function for polynomials.

In this paper, we investigate the frequency problem for the Rudin-Shapiro function along irreducible polynomials.
\bigskip

Let $q$ be an odd prime power and $\Fq$ be the finite field of $q$ elements. For a monic polynomial $f\in\Fq[t]$ of degree $n\geq 2$, define the \emph{Rudin-Shapiro function} $R(f)$ as
\begin{equation*}
R(f)=\sum_{i=1}^{n-1}f_if_{i-1} \quad \text{with } f=t^n+f_{n-1}t^{n-1}+\dots+f_0\in\Fq[t].
\end{equation*}

Let $\cP(n)$ be the set of \emph{monic irreducible} polynomials of degree $n$. It is well-known, that
\begin{equation}\label{eq:PNT}
\frac{q^n}{n}-2\frac{q^{n/2}}{n}\leq \#\cP(n)\leq \frac{q^n}{n}    
\end{equation}
(see for example \cite[Theorem~3.25]{LiNi97}).
For $\gamma\in\Fq$, let
$$
\cR_n(\gamma)=\{f\in\cP(n): R(f)=\gamma \}.
$$
Our goal is to prove that the size of $\cR_n(\gamma)$ is asymptotically the same for all $\gamma$.

Our main result is the following theorem.

\begin{theorem}\label{thm:1} 
Let $q$ be an odd prime power. 
There exists a $\delta>0$  such that for any $\gamma\in \Fq$ and $n\geq 3$ we have
$$
\#\cR_n(\gamma) = \frac{\#\cP(n)}{q} +O\left( n^{3/2}q^{(1-\delta)n +1}\right),
$$
where the implied constant  is absolute. One can choose $\delta =1/28$.
\end{theorem}

We remark, that  for $n=2$ the result follows from the Weil bound, see comment before \cite[Problem~8]{MeWi22}. 

By Theorem~\ref{thm:1}, we have that $$
\#\cR_n(\gamma) \sim \frac{\#\cP(n)}{q} 
$$
if either $n\rightarrow \infty$ or $q \rightarrow \infty$ provided $n>2/\delta$.

\section{Outline of the proof}


For a  positive integer $n$,  let $\cG(n)\subset \Fq[t]$ be the set of polynomials of degree $n$ and $\cM(n)\subset \cG(n)$ be the set of \emph{monic} polynomials of degree $n$.


For $\ell\geq 0$, write
$$
S^{(\ell)}(f)=\sum_{i=\ell}^nf_if_{i-\ell} \quad \text{with }  f=f_nt^n+f_{n-1}t^{n-1}+\dots+f_0\in\Fq[t].
$$
For $\ell=1$, we simply write
\begin{equation*}
S(f)=S^{(1)}(f).
\end{equation*}
Clearly, for monic polynomials $f$ we have 
\begin{equation}\label{eq:lin_red}
R(f)=S(f)-f_{n-1}.
\end{equation}
\bigskip

In order to prove Theorem~\ref{thm:1}, we introduce Vaughan's identity for polynomials in Section~\ref{sec:vau}. Then the main tool is to investigate single and double exponential sums in terms of $S(f)$ in Section~\ref{sec:qu}. The bounds of the exponential sum estimates are given in terms of the rank of quadratic forms which is studied in Sections \ref{sec:rev} and \ref{sec:rank}. Then we conclude the proof of Theorem~\ref{thm:1} in Section~\ref{sec:proof}.


\bigskip

For given functions $F$ and $G$, the notations $F\ll G$, $G \gg F$ and $F =O(G)$ are all equivalent to the statement that the inequality $|F| \leq c|G|$
holds with some constant $c > 0$. 

\section{preliminaries}
Let $\tau(f)$ denote the number of monic divisors of $f\in\Fq[t]$. First, we state the following bound on $\tau(f)$.

\begin{lemma}\label{lemma:tau}
For any $\varepsilon>0$ we have
$$
\tau(f) \ll \min\left\{ |f|^{(2+\varepsilon)/\log \deg f},2^{\deg f}
\right\},
$$
where the implied constant can depend on $\varepsilon$.
\end{lemma}
\begin{proof}
The first bound follows from \cite[Lemma~16]{Mer24+}, while the second bound follows from the fact that a polynomial $f$ has at most $\deg f$ irreducible divisor. 
\end{proof}

We also have the following bound on the second moment of $\tau$, see \cite[Lemma~8]{BiLe}.

\begin{lemma}\label{lemma:tau_moment}
    We have
    $$
\sum_{f\in \cM(n)}\tau(f)^2\leq 4n^3 q^n.
    $$
\end{lemma}

\section{Auxiliary results}
\subsection{Exponential sums of quadratic forms of polynomials}\label{sec:qu}

Let $Q$ be a quadratic form on $\Fq^n$. Then, $Q$ can be written as $Q(\bx)=\bx^TM\bx$ for some symmetric matrix $M\in\Fq^{n\times n}$. The \emph{rank} of $Q$ is the rank of the matrix $M$. The first result we need is the following bound. For completeness, we also include the proof.   

\begin{lemma}\label{lemma:gauss}
Let $q$ be an odd prime power and let $Q$ be a quadratic form, and let $L$ be a linear form on $\Fq^n$. Assume that $Q$ has rank at least $r$. Then
$$
\left|\sum_{\x\in \Fq^n}\psi(Q(\x)+L(\x))\right|\leq q^{n-r/2}
$$
for any non-trivial additive character $\psi$ of $\Fq$.
\end{lemma}

\begin{proof}
Write $Q(\x)=\x^TM\x$ for some symmetric matrix $M\in\Fq^{n\times n}$ of rank at least $r$ and let $B$ be the corresponding bilinear form 
$$
B(\x,\y)=\x^TM\y.
$$
Write also $B_\x(\y)=B(\x,\y)$. Let $K\leq \Fq^n$ the subspace of vectors $\x$ such that $B_\x=0$. It's codimension is the rank of $M$ and thus of the quadratic form $Q$.

We have
\begin{align*}
&\left|\sum_{\x\in \Fq^n}\psi(Q(\x)+L(\x))\right|^2\\
&=\sum_{\x,\h\in \Fq^n}\psi(Q(\x+\h)+L(\x+\h))\overline{\psi(Q(\x)+L(\x))}\\
&=\sum_{\h\in \Fq^n}\psi(Q(\h)+L(\h)) 
\sum_{\x\in \Fq^n}
\psi(2B_\h(\x)).
\end{align*}
If $\h\not \in K$, then $B_\h$ is a non-zero linear form and thus 
$$
\sum_{\x\in \Fq^n}
\psi(2B_h(\x))=\sum_{\x\in \Fq^n}
\psi(\x)=0.
$$
For $\h\in K$, we have
$$
\sum_{\x\in \Fq^n}
\psi(2B_\h(\x))=q^n.
$$
The assertion follows by noticing $\# K=q^{n-r}$.
\end{proof}

\subsection{Reversed polynomials}\label{sec:rev}
For a polynomial $a\in\Fq[t]$ of form $a=a_nt^n+\dots + a_1t+a_0$ let $a^*$ be the polynomial with reversed coefficients, that is,
$$
a^*(t)=t^na(t^{-1})=a_0t^n+\dots + a_{n-1}t+a_n\in\Fq[t].
$$
First we show, that the quantities $S^{(\ell)}(a)$ can be determined by using the reversed polynomial $a^*$ of $a$.

\begin{lemma}\label{lemma:star_coeff}
    For any polynomial $a\in\cG(n)$, we have 
    $$
    a^*\cdot a=\sum_{\ell=0}^n S^{(\ell)}(a)\left(t^{n+\ell}+t^{n-\ell}\right). 
    $$
\end{lemma}
\begin{proof}
The coefficient of the term $t^{n-\ell}$ in $a^*\cdot a$ is
$$
\sum_{ i+j=n-\ell} a_{n-i}a_{j}=\sum_{j=0}^{n-\ell} a_{j+\ell}a_{j}=S^{(\ell)}(a).
$$
Moreover, the coefficient of $t^{n-\ell}$ is the same as the coefficient of $t^{n+\ell}$.
\end{proof}

Next, we investigate the following type of quadratic equations in terms of the polynomials $a$ and its reversed $a^*$.
\begin{lemma}\label{lemma:quadratic_equation}
Let $\varepsilon >0$.
For given integer $n$, and polynomial $f\in \cG(2n)$, let $N(f)$ be the number of polynomials $a\in\cG(n)$ with 
\begin{equation}\label{eq:quadratic_equation}
a^*\cdot a=f.
\end{equation}
Then
$$
N(f)\ll \min\left\{ 2^n,q^{(2+\varepsilon)\frac{n}{\log_q n}}\right\},
$$
where the implied constant may depend on $\varepsilon$.
\end{lemma}

\begin{proof} Clearly, for any polynomials $a,b \in \Fq[t]$, we have
\begin{equation}\label{eq:star}
    (a\cdot b)^*=a^* \cdot b^*.
\end{equation}

Now assume that $f$ has the form $$
f=b^*\cdot b \quad \text{for some } b\in \cG(n),
$$
since otherwise the bound trivially holds.

If $a$ is a polynomial satisfying \eqref{eq:quadratic_equation}, put
$$
d=\gcd(a,b).
$$
Then by \eqref{eq:star}
$$
f= d \cdot \frac{b}{d} \cdot d^* \cdot \left(\frac{b}{d}\right)^*,
$$
and as
$$
\gcd(a/d,b/d)=1 \mand \gcd((a/d)^*,(b/d)^*)=1 
$$
we must have 
$$
a=d \cdot \left(\frac{b}{d}\right)^*. 
$$
That is, the divisor $d$ uniquely determines $a$. Whence, the number of polynomials $a$ with \eqref{eq:quadratic_equation} is at most $\tau(b)$. Then the bound follows from Lemma~\ref{lemma:tau}.
\end{proof}

 \subsection{Rank of quadratic forms}\label{sec:rank}
For a fixed $a\in \cM(k)$ the map $h\mapsto R(ah)$ with $h\in\cG(n-k)$ is a quadratic form in the coefficients of $h$. First we investigate its rank. 

\begin{lemma}\label{lemma:rank-g}
Let $a\in\cM(k)$ and for $h\in\cG(n-k)$ consider the map $Q_a: h\mapsto S(ah)$, which is a quadratic form in the coefficients of $h$. If $k<n/2$, then it has rank at least $\rk Q_a \geq n-k-1$.
\end{lemma}

\begin{proof}
Let $B$ be the symmetric bilinear form so that $S(h)=B(h,h)$. Specially,
\begin{equation}\label{eq:symm-bilin}
    B(t^i,t^j)=
\left\{
\begin{array}{cl}
  1/2   & \text{if } |i-j|=1, \\
  0   &  \text{otherwise},
\end{array}
\right. \quad 0\leq i,j\leq n.
\end{equation}

Put
$$
K=\{g\in\cG(n-k): h\mapsto B(ag,ah) \text{ is the zero map}\}
$$
which is a linear subspace of $\cG(n-k)$. Then 
\begin{equation}\label{eq:codim}
    \rk Q_a = n-k+1-\dim K. 
\end{equation}

Write 
$$
r^{(i)} \equiv t^i \mod a, \quad \deg r^{(i)}<\deg a
$$
and

$$
b^{(i)} = 
\left\{
\begin{array}{cl}
t^i     &  \text{if } i< k,\\
t^i -r^{(i)}     & \text{otherwise},
\end{array}
\right.
\quad i =0,1,\dots, n.
$$
Clearly, for each polynomial $f\in\cG(n)$ of form
$$
f=\sum_{i=0}^n f_i t^i
$$
one has
$$
f=\sum_{i=0}^n c_i b^{(i)} \quad \text{ for some } c_0,c_1,\dots,
$$
moreover
\begin{equation}\label{eq:divrep}
a\mid f \quad \Leftrightarrow \quad c_0=\dots =c_{k-1}=0, \text{ and }c_k=f_k,\dots, c_n=f_n.
\end{equation}
Indeed, if $a\mid f$, then 
\begin{align*}
    f= f-(f\bmod a)=\sum_{i=0}^n f_i (t^i-r^{(i)})=\sum_{i=k}^n f_i b^{(i)},
\end{align*}
as $t^i=r^{(i)}$ for $0\leq i<k$. Hence, $b^{(k)},\dots, b^{(n)}$ is a basis of the vector space of polynomials of form $ah$, $h\in \cG(n-k+1)$.

Now assume, that $g\in K$ and write
$$
ag=\sum_{i=0}^n d_i t^i.
$$
By \eqref{eq:divrep}, we can also write
\begin{equation}\label{eq:divrep2}
  ag=\sum_{i=k}^n d_i b^{(i)}.  
\end{equation}

As $g\in K$, we have by \eqref{eq:symm-bilin} that
for $\ell =k,\dots, n$
\begin{align}\label{eq:split}
0&=B(ag,b^{(\ell)})=B\left(\sum_{i=0}^n d_i t^i,t^\ell -r^{(\ell)}\right) \notag \\
&= B\left(\sum_{i=0}^n d_i t^i,t^\ell\right)-B\left(\sum_{i=0}^n d_i t^i,r^{(\ell)}\right) \\ \notag
&= \frac{d_{\ell+1}+d_{\ell-1}}{2}-B\left(\sum_{i=0}^n d_i t^i,r^{(\ell)}\right).
\end{align}
As $\deg r^{(\ell)}<k$, we have $B(t^i, r^{(\ell)})=0$ for $i>k$. The \eqref{eq:divrep2} and \eqref{eq:split} yield that
$$
d_{\ell+1}+d_{\ell-1}=2B\left(\sum_{i=0}^n d_i t^i,r^{(\ell)}\right)=2d_k B\left(  r^{(k)},r^{(\ell)}\right) , \quad \text{for } k\leq \ell<n.
$$
That  is, $d_k$ and $d_{k+1}$ determine all the values of $d_{k+2}, \dots, d_n$. Also, as 
$$
\sum_{i=0}^{k-1}d_it^i = \sum_{i=k}^{n} d_i (t^i \bmod a),
$$
the values $d_{0}, \dots, d_{k-1}$ are also uniquely determined. Thus
$\dim K=2$. Then the result follows from \eqref{eq:codim}.
\end{proof}

In the next lemma we investigate the rank of the quadratic form $h\mapsto S(ah)-S(bh)$ with $h\in\cG(n-k)$.

\begin{lemma}\label{lemma:num_of_solution_basic}
Let $a,b\in \cM(k)$ be polynomials such that
\begin{equation}\label{eq:star_assumption}
a^*\cdot a \neq b^*\cdot b.    
\end{equation}
Then, the quadratic form $h\mapsto S(ah)-S(bh)$ with $h\in\cG(n-k)$ has rank at least $n-2k-1$.
  \end{lemma}

\begin{proof}
Recall, that the symmetric bilinear form $B$ defined by \eqref{eq:symm-bilin} satisfies $S(h)=B(h,h)$. 

First consider the symmetric bilinear form $B_a$ defined by $$
B_a(h,g)=B(ah,ag),
$$ 
with $h,g\in\cG(n-k)$. Then
$$
B_a(t^i,t^j)=B(a\cdot t^i,a\cdot t^j)=\frac{1}{2} \left(S^{(|i-j-1|)}(a)+S^{(|j-i-1|)}(a) \right).
$$
Similarly, if $B_b$ is the symmetric bilinear form defined by $B_b(h,g)=B(bh,bg)$, $h,g\in\cG(n-k)$, we also have
$$
B_b(t^i,t^j)=B(b\cdot t^i,b\cdot t^j)=\frac{1}{2} \left(S^{(|i-j-1|)}(b)+S^{(|j-i-1|)}(b) \right).
$$
Whence, the symmetric bilinear form $B_{a,b}$ corresponding to $S(ah)-S(bh)$ satisfies
\begin{equation}\label{eq:Bab}
B_{a,b}(t^i,t^j)=\frac{1}{2} \left(S^{(|i-j-1|)}(a)+S^{(|j-i-1|)}(a)-S^{(|i-j-1|)}(b)-S^{(|j-i-1|)}(b) \right).    
\end{equation}

Let $M\in \Fq^{(n-k+1)\times (n-k+1)}$ be the symmetric matrix corresponding to $B_{a,b}$ with respect to the basis $1,t,t^2,\dots, t^{n-k}$. We show, that it has rank at least $n-2k-1$.  To do so
%
%
%
let
$\ell$ be the smallest integer such that $t^\ell \nmid a^*\cdot a - b^*\cdot b$. By \eqref{eq:star_assumption}, $\ell$ is well-defined. 
Also, as $a^*\cdot a - b^*\cdot b$ is symmetric, $\ell \leq k$. Then, by the definition of $\ell$ and by Lemma~\ref{lemma:star_coeff} we have
\begin{equation}\label{eq:S!=0}
S^{(\nu)}(a)=S^{(\nu)}(b) \quad \text{for } k-\ell < \nu \leq k \quad \text{and} \quad  S^{(k-\ell)}(a)\neq S^{(k-\ell)}(b).
\end{equation}
By definition of $S^{(\nu)}$, we also have, that 
\begin{equation}\label{eq:S=0}
    S^{(\nu)}(a)=S^{(\nu)}(b)=0 \quad \text{for }\nu >k.
\end{equation}

Then the submatrix
$$
(M_{i,j}) \quad \text{with } k-\ell +2\leq i\leq n-k+1 , \quad 1 \leq j\leq  n-2k+\ell
$$
of $M$ is an upper triangular matrix with nonzero elements in the diagonal by \eqref{eq:S!=0} and \eqref{eq:Bab}. Indeed, for $0\leq\nu,\mu<n-2k+\ell$ with $\nu>\mu$, we have
\begin{align*}
M_{k-\ell+2+\nu, 1+\mu}=
\frac{1}{2} \Big(&
S^{(|k-\ell +\nu-\mu |)}(a)+S^{(|\mu-\nu -k+\ell-2|)}(a)\\
& \quad -S^{(|k-\ell +\nu-\mu|)}(b)-S^{(|\mu-\nu -k+\ell-2|)}(b) 
\Big)=0.    
\end{align*}
by \eqref{eq:S=0}
as $|k-\ell +\nu-\mu|>k-\ell$ and $|\mu-\nu -k+\ell-2|>k-\ell +2$. Moreover, the values in the diagonal are  
\begin{align*}
&M_{k-\ell+2+\nu, 1+\nu}\\
&=
\frac{1}{2} \left(
S^{(|k-\ell|)}(a)+S^{(|-k+\ell-2|)}(a)-S^{(|k-\ell|)}(b)-S^{(|-k+\ell-2|)}(b) 
\right)\\
&=\frac{1}{2} \left(S^{(k-\ell)}(a)-S^{(k-\ell)}(b) \right)\neq 0    
\end{align*}
by \eqref{eq:S!=0} and \eqref{eq:S=0} as $|-k+\ell -2|>k-\ell$.

%
Thus it has rank $n-2k+\ell-1 $ and so
$$
\rk M\geq n-2k+\ell -1\geq n-2k-1.
$$
which proves the result.
\end{proof}










\subsection{Vaughan's identity for function fields}\label{sec:vau}
In this section we revise Vaughan's identity for function fields.
To do that, let us define the \emph{von Mangoldt function}
$\Lambda$ in $\F_q[t]$ by 
$$
\Lambda (f)=
\left\{
\begin{array}{cl}
 \deg P  & \text{if $f=P^r$ for some irreducible polynomial $P$;}  \\
  0   &  \text{otherwise.}
\end{array}
\right.
$$

Then we have the following result.

\begin{lemma}\label{lemma_Vau}
Let $\Psi:\F_q[t]\rightarrow \CC$ with $|\Psi(f)|\leq 1$. If $1\leq u,v\leq n$ two integers with $u+v<n$, then
$$
\sum_{f\in\cM(n)} \Lambda(f) \Psi(f)\ll n \Sigma_1 + n^{5/2} q^{n-(u+v)/2} \Sigma_2^{1/2}
$$
with
\begin{align*}
    \Sigma_1&=\sum_{|g|\leq q^{u+v}}\left|\sum_{|h|=q^{n}/|g|}\Psi(gh)\right|\\
    \Sigma_2&=\max_{v\leq i\leq n-u} \max_{|g_1|=q^{n-i}}\sum_{|g_2|=q^{n-i} }\left|\sum_{|h|=q^i}\Psi(hg_1)\overline{\Psi(hg_2)} \right|,
\end{align*}
where the summations are taken over \emph{monic} polynomials and the implied constant is absolute.
\end{lemma}

\begin{proof}
Let $\mu$ be the \emph{Möbius function} on $\F_q[t]$ defined by
$$
\mu (f)=
\left\{
\begin{array}{cl}
 (-1)^k  & \text{if $f=P_1\dots P_k$ for some distinct }  \\
   & \text{\quad irreducible polynomials $P_1,\dots, P_k$;}  \\
  0   &  \text{otherwise.}
\end{array}
\right.
$$
Then, one can write 
\begin{equation*}
    \Lambda(f)=\sum_{a\mid f}\mu(a)\deg (f/a),
\end{equation*}
where the sum is taken over only \emph{monic} polynomials. Then
\begin{equation*}
    \Lambda(f)= \sum_{\substack{a\mid f \\ |a|\leq q^u}} \mu(a) \deg (f/a) - \underset{\substack{ab\mid f\\ |a|\leq q^u, |b|\leq q^v}}{\sum  \ \sum } \mu(b)\Lambda(b) + \underset{\substack{ab\mid f\\ |a|> q^u, |b|>q^v}}{\sum  \ \sum } \mu(a)\Lambda(b).
\end{equation*}
It yields
\begin{equation*}
\sum_{f\in\cM(n)}\Lambda(f)\Psi(f)
= S_1-S_2+S_3,
\end{equation*}
where
\begin{align*}
S_1&=\underset{\substack{|ab|=q^n\\ |a|\leq q^u}}{\sum \ \sum}
 \mu(a) \deg (b) \Psi(ab),\\
 S_2&=\underset{\substack{|abc|=q^n\\ |a|\leq q^u, |b|\leq q^v}}{\sum  \ \sum \ \sum} \mu(a)\Lambda(b)\Psi(abc) ,\\
 S_3&=\underset{\substack{|abc|=q^n\\ |a|> q^u, |b|>q^v}}{\sum  \ \sum \sum } \mu(a)\Lambda(b)\Psi(abc).
\end{align*}
For $S_1$, we have
\begin{equation*}
S_1=\sum_{|a|\leq q^u}\mu(a)(n-\deg (a))\sum_{\substack{|b|=q^n/|a|}}\Psi(ab)\ll n\sum_{|a|\leq q^u}\left|\sum_{\substack{|b|=q^n/|b|}}\Psi(ab)\right|.
\end{equation*}
For $S_2$, observe, that 
$$
\sum_{b\mid z}\Lambda(b)=\deg (z).
$$
Then, we have
\begin{equation*}
\begin{split}
    S_2&=\sum_{r=n-u-v}^n \left(  \underset{\substack{|a|\leq q^u \, |b|\leq q^v\\ |ab|=q^{n-r}}}{\sum \ \sum}\mu(a)\Lambda(b)\right)\sum_{|c|=q^r}\Psi(abc)\\
    &\ll (u+v)\sum_{|z|\leq q^{u+v}}\left|\sum_{|c|=q^n/|z|}\Psi(zc)\right| .
\end{split}
\end{equation*}
For $S_3$, observe, that
$$
\sum_{a\mid c}\mu(a)=\left\{ 
\begin{array}{cl}
  1   & \text{if $c=1$}; \\
  0   & \text{otherwise}.
\end{array}
\right.
$$
Then, we have
\begin{equation*}
    S_3=\sum_{q^v\leq |b|\leq q^{n-u}}\Lambda(b)\sum_{|c|=q^n/|b|}\left(\sum_{\substack{a\mid c\\ |a|>q^u}}\mu(a)\right)\Psi(bc).
\end{equation*}
For $c\in\F_q[t]$, write
$$
\beta_c=\sum_{\substack{a\mid c\\ |a|> q^u}}\mu(a).
$$
Then, by the Cauchy-Schwarz inequality,
$$
\left|S_3\right|^2\leq \left(\sum_{q^v\leq |b|\leq q^{n-u}}\Lambda(b)^2\right)\sum_{q^v\leq |b|\leq q^{n-u}}\left|\sum_{|c|=q^n/|b|}\beta_c\Psi(bc)  \right|^2.
$$
For the first term, we have
$$
\sum_{q^v\leq |b|\leq q^{n-u}}\Lambda(b)^2 \leq n \sum_{ |b|\leq q^{n-u}}\Lambda(b)\leq n q^{n-u} .
$$
For the second term,
\begin{align*}
&\sum_{q^v\leq |b|\leq q^{n-u}}\left|\sum_{|c|=q^n/|b|}\beta_
c\Psi(bc)  \right|^2\\
&=\sum_{v\leq i\leq n-u} \sum_{|c_1|=q^{n-i}}\sum_{|c_2|=q^{n-i}} \beta_{c_1}\overline{\beta_{c_2}}\sum_{|b|=q^i}\Psi(bc_1)\overline{\Psi(bc_2)}\\
&\ll \sum_{v\leq i\leq n-u} \sum_{|c_1|=q^{n-i} } |\beta_{c_1}|^2 \sum_{|c_2|=q^{n-i} }\left|\sum_{|b|=q^i}\Psi(bc_1)\overline{\Psi(bc_2)} \right| \\
&\ll \sum_{v\leq i\leq n-u} \left(\sum_{|c_1|=q^{n-i} } |\beta_{c_1}|^2\right) \max_{|c_1|=q^{n-i}}\sum_{|c_2|=q^{n-i} }\left|\sum_{|b|=q^i}\Psi(bc_1)\overline{\Psi(bc_2)} \right|
\end{align*}
by $|\beta_{c_1}\overline{\beta_{c_2}}|\leq |\beta_{c_1}|^2/2 +|\beta_{c_2}|^2/2$ and by the symmetry.

By Lemma~\ref{lemma:tau_moment},
$$
\sum_{|c_1|=q^{n-i} } |\beta_{c_1}|^2\ll \sum_{|c_1|=q^{n-i} } |\tau(c_1)|^2 \ll n^{3}q^{n-i}.
$$

Thus
$$
|S_3|^2\ll n^5q^{2n-u-v} 
\max_{v\leq i\leq n-u}\max_{|c_1|=q^{n-i}}\sum_{|c_2|=q^{n-i} }\left|\sum_{|b|=q^i}\Psi(bc_1)\overline{\Psi(bc_2)} \right|.
$$
\end{proof}

\section{Proof of Theorem \ref{thm:1}}\label{sec:proof}
Let $\gamma\in \Fq$. By the orthogonality of additive characters, we have
\begin{equation*}
\cR(\gamma)=\frac{1}{q}\sum_{\psi}\sum_{f\in\cP(n)}\psi(R(f)-\gamma),
\end{equation*}
where the first summation is taken over all additive characters of $\Fq$. For the trivial additive character $\psi_0$, we have
\begin{equation*}
   \frac{1}{q}\sum_{f\in\cP(n)}\psi_0(R(f)-\gamma)=\frac{\#\cP(n)}{q}
\end{equation*}
which gives the main term.

The number of irreducible polynomials $f$ of degree $\deg f<n$ with $\deg f\mid n$ is $O(q^{n/2})$ by \eqref{eq:PNT}, thus for non-trivial additive characters $\psi\neq \psi_0$, we have
\begin{equation*}
  \begin{split}
    \sum_{f\in\cP(n)}\psi(R(f))
    &=\frac{1}{n}\sum_{f\in \cP(n)}\Lambda(f)\psi(R(f))\\
    &=\frac{1}{n}\sum_{f\in\cM(n)}\Lambda(f)\psi(R(f)) +O(q^{n/2}).
    \end{split}
\end{equation*}
Then,
by Lemma~\ref{lemma_Vau},
\begin{equation}\label{eq:Vau}
\sum_{f\in\cM(n)}\Lambda(f)\psi(R(f)) \ll  n \Sigma_1 + n^{5/2} q^{n-(u+v)/2} \Sigma_2^{1/2},
\end{equation}
where
$$
\Sigma_1=\sum_{|g|\leq q^{u+v}}\left|\sum_{|h|=q^{n}/|g|}\psi(R(gh))\right|
$$
and
$$
\Sigma_2=\max_{v\leq i\leq n-u} \max_{|g_1|=q^{n-i}}\sum_{|g_2|=q^{n-i} }\left|\sum_{|h|=q^i}\psi(R(hg_1))\overline{\psi(R(hg_2))} \right|
$$
with some $u,v$ to be chosen later.

In order to estimate $\Sigma_1$, observe, that 
by Lemma~\ref{lemma:rank-g}, the quadratic form $h\mapsto S(gh)$ with $\deg h\leq i$ has rank at least $n-\deg g-1$. If we consider it only on monic polynomials $h$, then its rank reduced by at most one and in this case $R(gh)=S(gh)+L_g(h)$ by \eqref{eq:lin_red} with some linear polynomial $L_g$ in terms of the coefficients of $h$.
Then, by Lemma~\ref{lemma:gauss},
\begin{align}\label{eq:Sigma1}
\Sigma_1&\ll \sum_{|g|\leq q^{u+v}} q^{n-\deg g- (n-\deg g-2)/2}\\ \notag
&=
\sum_{|g|\leq q^{u+v}} q^{(n-\deg g)/2+1}\\ \notag
&\ll q^{(n+2)/2} \sum_{i\leq u+v} q^{i/2}\\ \notag
&\ll q^{(n+u+v+2)/2}.
\end{align}

In order to estimate $\Sigma_2$, let us fix $v\leq i\leq n-u$ and consider
$$
I(g_1,g_1)=\sum_{|h|=q^i}\psi(R(g_1h)-R(g_2h)).
$$
with monic polynomials $g_1,g_2$ of degree $n-i$.
For given $g_1$, let 
$$
\cB(g_1)=\{g\in \cM(n-i): g^*g=g_1^*g_1 \}.
$$
If $g_1\not\in \cB(g_1)$, we have by Lemma~\ref{lemma:num_of_solution_basic}, that the rank of the quadratic form $h\mapsto S(g_1h)-S(g_2h)$ with $\deg h \leq i$  is at least $n-2(n-i)-1=2i-n-1$. If we consider it only on monic polynomials $h\in\cM(i)$, then its rank is reduced by at most one and in this case $R(g_1h)-R(g_2h)=
S(g_1h)-S(g_2h)+L_{g_1,g_2}(h)$ with some linear polynomial $L_{g_1,g_2}$ in the coefficients of $h$. Thus, in this case by Lemma~\ref{lemma:gauss} we get
$$
I(g_1,g_2)\ll q^{i-(2i-n-2)/2}=q^{n/2+1}.
$$
If $g_2\in\cB(g_1)$, we use the trivial bound
$$
I(g_1,g_2)\ll q^i.
$$

By Lemma~\ref{lemma:quadratic_equation} we have
$$
\#\cB(g_1)\ll \min\left\{ 2^n,q^{(2+\varepsilon)\frac{n}{\log_q n}}\right\}.
$$
By choosing $\varepsilon =1/2$, we get
$$
\#\cB(g_1)\ll  q^{n/14},
$$
where the implied constant is absolute. Then
\begin{equation}\label{eq:Sigma2}
\Sigma_2\ll \max_{v\leq i\leq n-u}\left(q^{n/14}q^i +q^{n-i}q^{n/2+1}  \right)\ll q^{15n/14-u}+q^{3n/2 -v+1}.
\end{equation}
Combining \eqref{eq:Vau}, \eqref{eq:Sigma1} and \eqref{eq:Sigma2}, we get
\begin{align*}
\sum_{f\in\cM(n)}\Lambda(f)\psi(R(f)) &\ll n q^{(n+u+v+2)/2} \\
&\quad + n^{5/2}q^{n-(u+v)/2}\left( q^{15n/28-u/2}+q^{3n/4-v/2+1/2}\right)\\
&\ll n q^{(n+u+v+2)/2} \\
&\quad + n^{5/2}\left( q^{43n/28-u-v/2  }+q^{7n/4-u/2-v+1/2}\right).
\end{align*}
Choosing
$$
u=\frac{3}{14}n \mand v=\frac{10}{14}n,
$$
we get
\begin{align*}
\sum_{f\in\cM(n)}\Lambda(f)\psi(R(f)) 
&\ll   n^{5/2}q^{27n/28 +1}.
\end{align*}

\section*{Acknowledgment}

The author wishes to thank Arne Winterhof for his valuable comments.

The author was was partially supported by NRDI (National Research Development and Innovation Office, Hungary) grant FK 142960 and by the János Bolyai Research Scholarship of the Hungarian Academy
of Sciences.


\end{document}